OpenStudy (anonymous):

Last month's electric bill was $125, while this month's is $100. This is a decrease of what percent?

OpenStudy (anonymous):

its simply 100-125/125 = -25/125 = -0.2 * 100 = 20% I think :D Hope its correct

OpenStudy (anonymous):

whats up Jose..